A box contains 35 coins in nickles and dimes. How many of each coin is there if the total of the two types of coins is $2.50? Please help me figure out the Let Statement.
Answer by stanbon(75887)

A box contains 35 coins in nickles and dimes. How many of each coin is there if the total of the two types of coins is $2.50? Please help me figure out the Let Statement.
------
Let the number of nickels be "n" and the number of dimes be "d".
----
Equations:
Quantity Eq.: n + d = 35
Value Eq.: 5n +10d= 250 cents
------------------------------------
Modify:
n + d = 35
n + 2d = 50
------
Subtract and solve for "d":
d = 15 (# of dimes)
---
solve for "n":
n + d = 35
n + 15 = 35
n = 20 (# of nickels)
=====================
Cheers,
Stan H.